a) Find an equation of the sphere with center (3, −6, 4) and radius 5.

b) Use an equation to describe its intersection with each of the coordinate planes. (If the sphere does not intersect with the plane, enter DNE.)

intersection with xy-plane    
intersection with xz-plane    
intersection with yz-plane
Answer #1

Genegal eqn of sphere @_gy?+Ly=b)+2=4572 - If ladius Ca, b, c) centre of sphere Centre (3, -6,4) & radilss (2=3){ [63?(2-4927